Solve the differential equations using integrating factor.

1. x dy/dx - y = x raise to power of 3 cosx, given that y=0 and x=π.

2. ( 1 + x square) dy/dx + 3xy = 5x given y=2 and x=1


1
Expert's answer
2021-02-01T08:23:58-0500
  1. In the first case we can even guess the integrating factor (as we can see the derivative of a quotient on the left). Let's calculate it directly anyway, first we will divide both sides by "x" to find "y' + (-\frac{1}{x}) y = x^2 \cos x" : "P(x)=-\frac{1}{x}", "I(x) = e^{\int P(t)dt} = \frac{1}{x}". By multiplying both sides by "I(x)" we find "\frac{y'}{x}-\frac{y}{x^2}=x\cos x". Now we have "(\frac{y}{x})' = x \cos x". By integrating both sides we find "\frac{y}{x} = x\sin x+\cos x + C" and thus "y(x) =x^2 \sin x + x\cos x + Cx". Let's find the constant : "0 = \pi^2 \cdot 0 + \pi \cdot (-1) + C \pi" and so "C=1", "y(x)=x^2 \sin x + x\cos x + x".
  2. Let's start by dividing both sides by "(1+x^2)" : "y'+\frac{3x}{1+x^2} y = \frac{5x}{1+x^2}". Thus "P(x) = \frac{3x}{1+x^2}", "I(x) = e^{\int P(t) dt} = e^{\frac{3}{2} \ln(1+x^2)}=(1+x^2)^{3/2}". Now by multyplying both sides by "I(x)" we find "y'(1+x^2)^{3/2} +3x\cdot (1+x^2)^{1/2}y=5x\cdot (1+x^2)^{1/2}", "((1+x^2)^{3/2}y)' = 5x \cdot (1+x^2)^{1/2}". Now by integrating both sides we find "(1+x^2)^{3/2} y = \frac{5}{2} \cdot \frac{2}{3} (1+x^2)^{3/2} + C" and so "y(x)=\frac{5}{3} + \frac{C}{(1+x^2)^{3/2}}", now let's find the constant : "y(1) = \frac{5}{3} + \frac{C}{2^{3/2}} = 2", "C = \frac{2\sqrt{2}}{3}" and thus the solution is "y(x) = \frac{5}{3} + \frac{2\sqrt{2}}{3(1+x^2)^{3/2}}".